Solve
px^2+2qxy-pq-2xz=0
1
Expert's answer
2020-11-03T17:01:42-0500
"f(x, y,z,p,q)=px^2+2qxy-pq-2xz=0"

We have the system of Charpit’s auxiliary equations as


"\dfrac{dx}{-f_p}=\dfrac{dy}{-f_q}=\dfrac{dz}{-pf_p-qf_q}=\dfrac{dp}{f_x+pf_z}=\dfrac{dq}{f_y+qf_z}"


"\dfrac{dx}{q-x^2}=\dfrac{dy}{p-2xy}=\dfrac{dz}{-px^2+pq-2qxy+pq}="

"=\dfrac{dp}{2px+2qy-2z-2px}=\dfrac{dq}{2qx-2qx}"

Now "dq=0" or "q=a"


"px^2+2axy-pa-2xz=0"

"p(x^2-a)=2x(z-ay)"

"p=\dfrac{2x(z-ay)}{x^2-a}"

"dz=pdx+qdy"

"dz=\dfrac{2x(z-ay)}{x^2-a}dx+ady"

"\dfrac{dz-ady}{z-ay}=\dfrac{2x}{x^2-a}dx"

Integrate both sides


"\ln(|z-ay|)=\ln(|x^2-a|)+\ln c"

"z=c(x^2-a)+ay"